A blimp is cruising slowly at low altitude, filled as usual with helium gas. Its maximum useful payload, including crew and cargo, is 1.52E+3 kg. The volume of the helium-filled interior space is 5.51E+3 m3. The density of helium gas is 0.16 kg/m3, and the density of hydrogen is 0.081 kg/m3. How much more payload could the blimp carry if you replaced the helium with hydrogen?
Answer #1

given

m_crewandcargo = 1.52*10^3 kg
V_He = 5.51*10^3 m^3
rho_He = 0.16 kg/m^3
rho_H = 0.081 kg/m^3


if the balloon is filled with He,
m_He = rho_He*V_balloon

= 0.16*5.51*10^3

= 881.6 N

if the balloon is filled with H
m_He = rho_He*V_Balloon

= 0.081*5.51*10^3

= 446.31 N


the amount more payload could the blimp carry = 881.6 - 446.3

= 435 kg <<<<<<<<<<-------------------Answer